OpenStudy (anonymous):

For which operations is the set {–1, 1} closed? Choose all answers that are correct. A. addition B. subtraction C. multiplication D. division

OpenStudy (anonymous):

@ganeshie8 @geerky42 @zzr0ck3r @tommynaiter

OpenStudy (zzr0ck3r):

is 1+1 in the set?

OpenStudy (zzr0ck3r):

is 1-1 in the set?

OpenStudy (zzr0ck3r):

is -1*-1=1, -1*1=-1 all in the set?

OpenStudy (anonymous):

I'm so confused

OpenStudy (zzr0ck3r):

is -1/1 and 1/1 and -1/-1 and 1/-1 in the set?

OpenStudy (anonymous):

What does ion the set mean? what makes something in the set?

OpenStudy (zzr0ck3r):

a set is closed wrt operation if for any 2 elements in the set, the result of the operation on those 2 elements is in the set

OpenStudy (zzr0ck3r):

since 1 and 1 are both in the set, and 1+1 = 2 is not. The set is NOT closed with respect to addition

OpenStudy (zzr0ck3r):

similarly 1-1 = 0 is not in the set, so it is not closed with respect to subtraction.

OpenStudy (zzr0ck3r):

what about multiplication?

OpenStudy (anonymous):

ohh I get it

OpenStudy (anonymous):

So it's not closed with subtraction

OpenStudy (anonymous):

or addition

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

but it is for multipplication

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

And division

OpenStudy (zzr0ck3r):

right, does it still work for division if we throw in 0?

OpenStudy (zzr0ck3r):

or any other number?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

?

OpenStudy (zzr0ck3r):

correct, say we throw in 3, then we need 1/3 and then we need (1/3)/3 = 1/9...

OpenStudy (anonymous):

Thank you! were so helpful!

OpenStudy (zzr0ck3r):

np